Samson is 6 years old. His mother is 30 years old. How old will Samson be when his mother is exactly 4 times as old as he is?
Answer by Maths68(1474)

Samson is 6 years old. His mother is 30 years old. How old will Samson be when his mother is exactly 4 times as old as he is?
Present age of mother = 30
Present age of son = 6
Let
mother will be 4 times in = x years
After x years
Mother will be = 30+x
Son will be = 6+x
Then
4 * (son's age) = Mother's age
4*(6+x)=30+x
24+4x=30+x
4x-x=30-24
3x=6
3x/3=6/3
x=2
After 2 years mother's age will be 4 times of son's age.

Check
=====
After 2 years
Son will be = 6+x = 6+2 = 8 years old
Mother will be = 30+x = 30+2= 32 years old (which is 4 times of 8)